OpenStudy (anonymous):

Identify the slope y =-3/4x-75

OpenStudy (skullpatrol):

In the equation y=mx+b, m is called the slope.

OpenStudy (anonymous):

no thats the equation not the slope

OpenStudy (anonymous):

Slope= y=mx+b M is the slope, B is the Y axis. In your equation, -3/4 is your slope and 75 is your y axis

OpenStudy (anonymous):

3/4 is the slope

OpenStudy (skullpatrol):

$$\Huge y =\color{red}{\frac{-3}{4}} x-75$$ $$\Huge y=\color{red}mx+b$$

OpenStudy (anonymous):

^